OpenStudy (anonymous):

A table for y=x^2-64 is given below. Solve each equation or inequality. b. X^2-64<0

OpenStudy (anonymous):

your quadratic has zeros at -8 and 8, and since it is a parabola that faces up, it is negative between the zeros and positive outside of them (draw a quick sketch) so your answer is \[-8<x<8\]

OpenStudy (anonymous):

Okay so for x^2-64=0 it would be -8,8?

OpenStudy (anonymous):

the question asks for a solution to an inequality, your answer must be an interval

OpenStudy (anonymous):

The only options for that answer would be a.-8,8 b -8,0 e. -9,9 g. -64,0

OpenStudy (anonymous):

I thought that since it was set to 0 already It would be -8,8

OpenStudy (anonymous):

it is not set equal zero in your question. it asks where it is LESS THAN zero which is a synonym for "below the \(x\) axis

OpenStudy (anonymous):

Sorry, more options a. -8,8 b. 0,9 c. -8,0 d. -64. e. -9,9 f. 0 g. -64,0 and h. no real solution

OpenStudy (anonymous):

if they wrote the answer in "interval notation" it would be \((-8,8)\)

OpenStudy (anonymous):

means the same thing as \(-8<x<8\)

OpenStudy (anonymous):

okay, that makes since for the question of x^2-64>0

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

you wrote in the question \[x^2-64<0 \]

OpenStudy (anonymous):

is it less than \(x^2-64<0\) or greater than \(x^2-64>0\)

OpenStudy (anonymous):

it's a three part question.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the last part is x^2-64>0

OpenStudy (anonymous):

if you know where it is less than zero, namely on \((-8,8)\) then you know it is greater than zero everywhere else, namely \((-\infty,-8)\cup (8,\infty)\)

OpenStudy (anonymous):

okay, But does that mean that it's x<-8 and x>8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well, actually it means \(x<-8\) OR \(x>8\)

OpenStudy (anonymous):

there is no such number that is both, it is an OR statement, not an AND statement

OpenStudy (anonymous):

Okay, is there a way I can run my questions by to check the answers, some of them I am unsure that they are correct?

OpenStudy (anonymous):

you want me to look?

OpenStudy (anonymous):

that would be awesome!

OpenStudy (anonymous):

sure, i am good for another 5 minutes or so shoot

OpenStudy (anonymous):

Find the maximum y-value on the graph of y=f(x).. f(x) -x^2+4x+9 max is 13?

OpenStudy (anonymous):

To solve by completing the square, what value should you add to each side. x^2+10x=-4. (25)

OpenStudy (anonymous):

first one is right

OpenStudy (anonymous):

second one is right

OpenStudy (anonymous):

Use the quadratic formula to find any x-intercepts on the graph of the equation. y=x^2+6x-10 x= -3+sqrt19, -3-sqrt19

OpenStudy (anonymous):

third one is right

OpenStudy (anonymous):

for y=x^2-5 do the following identify the vertex and compare the graph of y=f(x) to the graph of y=x^2 state any transformations used vertex is (0,-5) and y=x^2 shifted 5 units downwards

OpenStudy (anonymous):

that is right too

OpenStudy (anonymous):

and my graph looks like this